OpenStudy (anonymous):

x+y=4 2x-y=5 What is the solution to this system?

OpenStudy (anonymous):

use elimination method

OpenStudy (anonymous):

add down using elimination method

OpenStudy (anonymous):

add solve for x then solve for y I think it will be easier

OpenStudy (anonymous):

multiply both side of x+y=4 by 5 and multiply both side of 2x-y=5 by 4

OpenStudy (anonymous):

I think adding two equation will give you in x terms only... no ned to go in a roundabout way

OpenStudy (anonymous):

Thank you SO MUCH! all of you @gerryliyana @Rosh007 @Mr.Me I appreciate 100% :p

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

welcome..so u got it... O glad..

OpenStudy (anonymous):

Is the answer suppose to be a decimal tho??

OpenStudy (anonymous):

because i got -0.3333333333.. etc.

OpenStudy (anonymous):

what did u get it??

OpenStudy (anonymous):

show in -1/3

OpenStudy (anonymous):

by using the elimination method ... Oh okay @gerryliyana then plug that in to x in x+y=4 ?

OpenStudy (anonymous):

yes... first you nd to eliminate one variable... sinc there are two equations and two unknowns.. it would is possible... but you won't get -0.3333.. in this case

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

NO!!! you have two use the equations which are available to you

OpenStudy (anonymous):

What would it be @Rosh007 because thats what i got, am i doing it wrong then??

OpenStudy (anonymous):

ur right @JessNicole

OpenStudy (anonymous):

Try using substitution and see if you got the same as me. (:

OpenStudy (anonymous):

@jazy you made a mistake here Solve for x in the first equation. x + y = 4 Subtract y. x = 4 - y Substitute the value of x into the second equation. Solve for y. 2(4 - y) = 5 ----- mistake

OpenStudy (anonymous):

be careful @jazy 2(4 - y) = 5 ???? the right is 2(4 - y) - y= 5

OpenStudy (anonymous):

Oh okay now I see.

OpenStudy (anonymous):

@jazy THANK YOU!!! it makes sence but @Rosh007 how is it a mistake?

OpenStudy (anonymous):

be careful @jazy the mistake at 2(4 - y) = 5 ???? the right is 2(4 - y) - y= 5

OpenStudy (anonymous):

@JessNicole please look gerryliyana's reply

OpenStudy (anonymous):

Ohh okayy, i see now soo, it'd be Distribute 8-y=5?? @Rosh007 and @jazy ??

OpenStudy (anonymous):

jazzy missed one y there

OpenStudy (anonymous):

@gerryliyana Thank youuuu!!:) for catching that mistake, i fixed it on my paper!

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I prefer substitution. x + y = 4 2x - y = 5 Solve for x in the first equation. x + y = 4 Subtract y. x = 4 - y Substitute the value of x into the second equation. Solve for y. 2(4 - y) - y = 5 Distribute. 8 - 2y - y= 5 Subtract 8. -2y - y = -3 -3y = -3 Divide both sides by -3. y = 1 Now plug in the value of y(1) into any equation to solve for x. x + y = 4 x + 1 = 4 Subtract 1 x = 3 ---------------------------------- So x = 3 and y = 1

OpenStudy (anonymous):

@gerryliyana What would it be then?

OpenStudy (anonymous):

2(4 - y) - y= 5 8 - 3y = 5

OpenStudy (anonymous):

2(4 - y) - y= 5 8 - 3y = 5

OpenStudy (anonymous):

Thank you catching that mistake! (:

OpenStudy (anonymous):

It should be right now!

OpenStudy (anonymous):

yes you got it jazzy.. but a very simpl method is there

OpenStudy (anonymous):

@JessNicole 2(4 - y) - y= 5 8 - 3y = 5

OpenStudy (anonymous):

@jazy @Rosh007 @gerryliyana THANK YOUUUUUUUUU!! SO MUCH EVERYONE!!:D .. I'd give all of you guys each a medal if i could!!

OpenStudy (anonymous):

sice answer is already wrote here( which I don't like to do write directly) it is for you to know that there is a simple method than this

OpenStudy (anonymous):

@jazy you are welcome

OpenStudy (anonymous):

@JessNicole ur welcome..., happy to help u :)

OpenStudy (anonymous):

@Rosh007 There's an easier way ? cause you can teach it to me ..this stuff is hard for me, i'm always confused.

OpenStudy (anonymous):

actually it is derivd from the method mentioned here.. It is actually a special case of this method.. it is applicabl here since in one equation +y and in other equation -y is there actually the problem is simple x+y=4 2x-y=5  so when we add both equation (LHS with LHS and RHS with RHS) you get 3x =9 which can be simplified quickly and reach the answer fast